Ampere’s law assignment
Two long parallel wires are 10 cm apart and carry currents of 6 A and 4 A. what is the field B at 4 A carrying wire due to 6 A wire.
1
Expert's answer
2020-05-15T16:10:54-0400

Explanations & calculations


Consider the sketch attached.


  • Consider of the magnetic field induced on the point P by AB finite length conductor.
  • In this question Biot - Savart law is applied which is

"\qquad\qquad
dB = \frac{\mu_0i}{4\pi}\bigg(\frac{ldl*\sin\theta}{r^2}\bigg) \cdots(1)" ; symbols have the standard meanings


Proof:

  • In the equation above the integral is ldl & have 3 variables within the brackets, which should be reduced to evaluate (lets turn it into an integral of d"\small \theta" )
  • "\qquad\qquad
\begin{aligned}
\small r &= \small \frac{d}{\sin\theta_1}\cdots(2)\\
\small l &= \small \frac{d}{\tan\theta_1} = d\cot\theta\cdots(3)\\

\end{aligned}"
  • Now find dl with d"\small \theta" by differentiating (3) w.r.t "\theta"

"\qquad\qquad\qquad
\begin{aligned}
\small dl &= \small -\frac{d}{\sin^2\theta}*d\theta
\end{aligned}"

  • Now re-writing equation (1) we get,

"\qquad\qquad
\begin{aligned}
\small dB &= \small \frac{\mu_0i}{4\pi}\bigg(\frac{(\frac{-d}{\sin^2\theta_1}*d\theta)\sin \theta }{(\frac{d}{\sin\theta_1})^2}\bigg)\\
&= \small \frac{\mu_0i}{4\pi}\bigg(\frac{-\sin\theta_1}{d}d\theta\bigg)\\
&= \small \frac{-\mu_0i}{4\pi d}\big(\sin\theta d\theta\big)
\end{aligned}"

  • Now integrate over the AB length of the conductor to obtain the net magnetic field on P

"\qquad\qquad
\begin{aligned}
\small \intop dB &= \small \frac{-\mu_0i}{4\pi d} \intop_{\theta_1}^{(\pi-\theta_2)} \sin\theta d\theta\\
\small B &= \small \frac{-\mu_0i}{4\pi d}\big(-\cos\theta_2 - \cos\theta_1\big)\\
&= \small \frac{\mu_0i}{4\pi d}\big(\cos\theta_2 + \cos\theta_1\big)
\end{aligned}"

  • Now this is for the effect from a finite length conductor & as the conductor is infinite, "\theta_1,\theta_2 \to0 \,\,\text{and}\,\, \cos\theta_1,\cos\theta_2\to 1"


  • Therefore, magnetic field strength on a point due to a conductor of infinite length becomes,

"\qquad\qquad\qquad
\begin{aligned}
\small B = \small \frac{\mu_0i}{2\pi d}
\end {aligned}"


Answer

  • Magnetic field on 4A carrying conductor due to the other conductor,

"\qquad\qquad
\begin{aligned}
\small B &= \small \frac{4\pi\times10^{-7}TmA^{-1}\times6A}{2\pi\times0.1m}\\
&= \small \bold{1.2\times10^{-5}T}
\end{aligned}"